OpenStudy (anonymous):

Does anyone know how to do this math question::: Joshua Bennett purchased a television for $245.95 with 40.25% down. He will make 6 payments of $25.83 each. What is the total amount paid (will include down payment) and finance charge in dollars?

OpenStudy (anonymous):

Help!!!11

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}\%\ of \ b=\cfrac{{\color{red}{ a}}}{100}\times b\\ \quad \\ {\color{red}{ 40.25}}\%\ of \ 245.95=\cfrac{{\color{red}{ 40.25}}}{100}\times 245.95\implies \square \\ \quad \\ \textit{plus 6 payments of 25.83}\implies 6\cdot 25.83 \\ \quad \\ \textit{total amount paid}=\square +(6\cdot 25.83)\)